OpenStudy (anonymous):

determine this whether is converge or diverge, if diverge, find its sum. (question going to be added now)

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} 2/(n^2+4n+3) \]

OpenStudy (anonymous):

sorry , if converge, find its sum

OpenStudy (anonymous):

Well, it converges, and you can compare with 2/n^2, which is a p-series with p > 1. But calculating the sum precisely is another thing...

OpenStudy (anonymous):

it converges by nth term test it goes to zero

OpenStudy (anonymous):

we use the geometry rule?

OpenStudy (anonymous):

for sum yes i guess we use geometric rule

OpenStudy (anonymous):

What geometric rule? This doesn't look like a geometric series, does it?

OpenStudy (anonymous):

sorry is series nor rule mistake typying

OpenStudy (anonymous):

if u do show then show the procedure that was is asking in the texbook

OpenStudy (anonymous):

i guess that is one way to find the sum if the result converge

OpenStudy (anonymous):

If it's a geometric sequence, what is r here? I did calculate some terms and they seem to approach 5/6, but I am somewhat numbed, I guess. Can't see how to find the precise sum of that.

OpenStudy (anonymous):

I mean, their sum seems to approach 5/6.

OpenStudy (anonymous):

if opositive then is diverger when is infinity

OpenStudy (anonymous):

thank so much for the help bmp

OpenStudy (anonymous):

Was that right? :-O

OpenStudy (anonymous):

the series can become : \[\sum_{n=1}^{\infty} {(\frac{1}{n+1}-\frac{1}{n+3})}=\sum_{n=1}^{\infty} {\frac{1}{n+1}}-\sum_{n=1}^{\infty} {\frac{1}{n+3}}\] We can see that \[\sum_{n=1}^{\infty} {\frac{1}{n+1}}=\frac{1}{2}+\frac{1}{3}+\frac{1}{4}+\frac{1}{5}+....\] \[\sum_{n=1}^{\infty} {\frac{1}{n+3}}=\frac{1}{4}+\frac{1}{5}+\frac{1}{6}+\frac{1}{7}+....\]So when we use subtraction here for \[n\to \infty\] Its sum is only \[\frac{1}{2}+\frac{1}{3}=\frac{5}{6}\] as bmp says

OpenStudy (anonymous):

Ahh, that makes sense. Kudos for seeing more than I did :-)